# Retrieval-Augmented Dynamic Prompt Tuning for Incomplete Multimodal Learning

### v1

@@ Abstract
Multimodal learning with incomplete modality is practical and challenging. Recently, researchers have focused on enhancing the robustness of pre-trained MultiModal Transformers (MMTs) under missing modality conditions by applying learnable prompts. However, these prompt-based methods face several limitations: (1) incomplete modalities provide restricted modal cues for task-specific inference, (2) dummy imputation for missing content causes information loss and introduces noise, and (3) static prompts are instance-agnostic, offering limited knowledge for instances with various missing conditions. To address these issues, we propose RAGPT, a novel Retrieval-AuGmented dynamic Prompt Tuning framework. RAGPT comprises three modules: (I) the multi-channel retriever, which identifies similar instances through a within-modality retrieval strategy, (II) the missing modality generator, which recovers missing information using retrieved contexts, and (III) the context-aware prompter, which captures contextual knowledge from relevant instances and generates dynamic prompts to largely enhance the MMT’s robustness. Extensive experiments conducted on three real-world datasets show that RAGPT consistently outperforms all competitive baselines in handling incomplete modality problems. The code of our work and prompt-based baselines is available at https://github.com/Jian-Lang/RAGPT.
@@ Introduction
$\bullet$ We conduct extensive experiments on three real-world datasets to evaluate RAGPT in comparison with 9 competitive baselines and the results confirm RAGPT’s effectiveness in addressing missing-modality issues.
@@
Missing-Modality in Multimodal Learning. Researchers have developed various robust methods for incomplete multimodal learning, which can be divided into three groups: (1) Joint learning methods (Zhao, Li, and Jin 2021; Wang et al. 2023; Yao et al. 2024) focus on distilling complex correlations from complete modalities to tackle missing-modality data. However, these methods require filling modality-incomplete inputs with masking values, which may cause unexpected behavior and introduce additional noise in the inference process. (2) Cross-modal generation methods (Lee et al. 2019; Yuan et al. 2021) primarily reconstruct the missing content by using remaining modalities. Researchers (Ma et al. 2021; Woo et al. 2023) directly employ VAE to generate the missing-modality based only on available modalities. Consequently, these methods inevitably face modality heterogeneity problems. (3) Prompt-based methods (Lee et al. 2023; Jang, Wang, and Kim 2024) represent a recent trend in this field, which introduces learnable prompts to help pre-trained MMTs address incomplete modalities.
@@
Setting of Missing Pattern We define the missing rate $\eta\%$ as the proportion of modality-incomplete data relative to the entire dataset. For each dataset, there are three possible cases of missing-modality: text missing, image missing, and both modalities missing. Text/image missing with a missing rate of $\eta\%$ indicates that there are $\eta\%$ instances consisting of texts/images and (1-$\eta$)instances that contain both modalities. Missing both modalities with a missing rate of $\eta\%$ indicates that there are $\frac{\eta}{2}\%$ instances consisting solely of images, $\frac{\eta}{2}\%$ instances consisting solely of text, and (1-$\eta$)instances that are complete, containing both modalities.
Implementation Details Following prior works (Lee et al. 2023; Jang, Wang, and Kim 2024), we utilize the pre-trained ViLT (Kim, Son, and Kim 2021) as our MMT backbone. The memory $\mathcal{B}$ for each dataset is constructed with the corresponding training set. The length $l$ of context-aware prompts is set to 2, the number of retrieved instances $K$ is chosen from $\{1,3,5,7,9\}$, and the prompt insertion layer $b$ is set to 2. We utilize the AdamW optimizer (Loshchilov and Hutter 2017) with a learning rate of $1\times 10^{-3}$ and total 20 epochs for optimizing the parameters. All experiments are conducted with an NVIDIA RTX 3090 GPU.
@@
To verify the superiority of RAGPT, we compare it with 9 competitive baselines on three datasets under a missing rate of $\eta\%=70\%$. From these results, we have the following observations:

### v2

@@ Abstract
Multimodal learning with incomplete modality is practical and challenging. Recently, researchers have focused on enhancing the robustness of pre-trained MultiModal Transformers (MMTs) under missing modality conditions by applying learnable prompts. However, these prompt-based methods face several limitations: (1) incomplete modalities provide restricted modal cues for task-specific inference, (2) dummy imputation for missing content causes information loss and introduces noise, and (3) static prompts are instance-agnostic, offering limited knowledge for instances with various missing conditions. To address these issues, we propose RAGPT, a novel Retrieval-AuGmented dynamic Prompt Tuning framework. RAGPT comprises three modules: (I) the multi-channel retriever, which identifies similar instances through a within-modality retrieval strategy, (II) the missing modality generator, which recovers missing information using retrieved contexts, and (III) the context-aware prompter, which captures contextual knowledge from relevant instances and generates dynamic prompts to largely enhance the MMT’s robustness. Extensive experiments conducted on three real-world datasets show that RAGPT consistently outperforms all competitive baselines in handling incomplete modality problems.
@@ Introduction
$\bullet$ We conduct extensive experiments on three real-world datasets to evaluate RAGPT in comparison with 9 competitive baselines and the results confirm RAGPT’s effectiveness in addressing missing-modality issues. The code of our work and prompt-based baselines is available at https://github.com/Jian-Lang/RAGPT.
@@
Incomplete Multimodal Learning. Researchers have developed various methods for incomplete multimodal learning, which can be divided into three groups: (1) Joint learning methods (Zhao, Li, and Jin 2021; Wang et al. 2023; Yao et al. 2024) focus on distilling complex correlations from complete modalities to tackle missing-modality data. However, these methods require filling modality-incomplete inputs with masking values, which may cause unexpected behavior and introduce additional noise. (2) Cross-modal generation methods (Lee et al. 2019; Yuan et al. 2021) primarily reconstruct the missing content by using remaining modalities. Researchers (Ma et al. 2021; Woo et al. 2023) directly employ VAE to generate the missing-modality based only on available modalities. Consequently, these methods inevitably face modality heterogeneity problems. (3) Prompt-based methods (Lee et al. 2023; Jang, Wang, and Kim 2024) introduces learnable prompts to help pre-trained MMTs address incomplete modalities.
@@
Setting of Missing Pattern We assume training set is fully available and define the missing rate $\eta\%$ as the rate of modality-incomplete samples in the test set: (1) text/image missing with $\eta\%$ indicates that there are $\eta\%$ image-only/text-only instances and (1-$\eta\%$) modality-complete instances. (2) both modalities missing with $\eta\%$ indicates that there are $\frac{\eta}{2}\%$ text-only instances, $\frac{\eta}{2}\%$ image-only instances and (1-$\eta\%$) modality-complete instances. We set missing rate $\eta\%$ = 70% by default. For training of each model, we simulate the same 70% missing rate to align model optimization well with test conditions, but allow each model to access the full modality information in the training set.
Implementation Details Following prior works (Lee et al. 2023; Jang, Wang, and Kim 2024), we utilize the pre-trained ViLT (Kim, Son, and Kim 2021) as our MMT backbone. The memory $\mathcal{B}$ for each dataset is constructed with the corresponding training and validation set without test data leakage. The length $l$ of prompts is set to 2, the number of retrieved instances $K$ is chosen from $\{1,3,5,7,9\}$, and the prompt insertion layer $b$ is set to 2. We utilize the AdamW optimizer with a learning rate of $1\times 10^{-3}$ and total 20 epochs for optimizing the parameters. All experiments are conducted with an NVIDIA RTX 3090 GPU.
@@
To verify the superiority of RAGPT, we compare it with 9 competitive baselines on three datasets under a missing rate of $\eta\%=70\%$. We have the following observations: